# Quantitative Introspection in Language Models: Tracking Internal States Across Conversation

### 19 Mar 2026 v1

@@ 3.3 Concept probes
We train probes for four emotive concept pairs: sad vs. happy (wellbeing), bored vs. interested (interest), distracted vs. focused (focus), and impulsive vs. planning (impulsivity). Each probe is a contrastive mean-difference direction (Zou et al., 2023): for a set of N=2024N=20-24 neutral questions (e.g., “Explain how photosynthesis works”), we generate completions under two opposing system prompts that induce the positive and negative poles of the concept, extract hidden-state activations from all layers, and compute the concept vector at each layer $l$ as:
@@
where $\bar{\mathbf{h}}_{l}^{+}$ and $\bar{\mathbf{h}}_{l}^{-}$ are the mean hidden-state representations (averaged over all assistant tokens) of completions under the positive and negative system prompts, respectively. We select the best layer $l^{*}$ as the one maximizing Cohen’s $d$ on a set of held-out evaluation texts (separate from the training prompts), with the search restricted to the middle 60% of layers, to lower the chance of selecting layers that mainly capture syntactic or grammatical differences between prompt and/or response structure. The probe score window then includes $l^{*}$ and its two nearest neighbours on each side.
Two of the four probes (wellbeing and impulsivity) were deliberately trained with opposite polarity relative to the self-report scale: the positive training pole is sad for wellbeing and planning for impulsivity, and in both cases it corresponds to the low end of the later self-report question. However, scores for these probes are sign-corrected before analysis so that higher probe scores align with higher self-report values in all results showed throughout this study. This design choice serves as a control, verifying that positive results are not an artifact of shared polarity conventions. Random-direction controls, using vectors of equal per-layer norm drawn from an isotropic distribution, serve as the primary null comparison.
Full details on system prompts, training questions, evaluation texts, and sign conventions are provided in Appendix B.
@@
Figure 1: Linear probes recover four interpretable internal directions in LLaMA-3.2-3B-Instruct. Panels A, C, E, and G show layer-wise Cohen’s $d$ sweeps for the sad-vs-happy (wellbeing), bored-vs-interested (interest), distracted-vs-focused (focus), and planning-vs-impulsive (impulsivity) probes; dashed lines mark the selected layers and the gray band marks the searched layer range. The best layers are 16, 14, 10, and 13, with peak $d$ values 3.34, 1.67, 1.99, and 3.60. For wellbeing and impulsivity, scores were sign-corrected so that larger values align with larger self-reported values in later experiments. Panels B, D, F, and H show the selected-layer score distributions on held-out evaluation texts; boxplots show text-level scores. The two poles separate in all four cases (Welch’s $t$-test: wellbeing, $d=3.34$, $p=7.21\times 10^{-13}$; interest, $d=1.67$, $p=9.45\times 10^{-6}$; focus, $d=1.99$, $p=5.73\times 10^{-7}$; impulsivity, $d=3.60$, $p=3.58\times 10^{-13}$), and all four survive BH correction across concepts.
@@
Figure 1 presents the probe-quality results for all four concepts in LLaMA-3.2-3B-Instruct. Panels A, C, E, and G show the layer-wise Cohen’s $d$ sweep for each probe; the selected best layers are 16, 14, 10, and 13, for wellbeing, interest, focus, and impulsivity respectively. Panels B, D, F, and H show the score distributions on held-out evaluation texts at the selected layer. The two poles separate clearly in all four cases (Welch’s $t$-test: all $p<10^{-5}$; Cohen’s d=1.67d=1.673.603.60), confirming that the trained directions capture meaningful structure. These probes serve as the internal-state readouts for all subsequent experiments.
@@ 3.4 Activation steering
Self-steering uses the same concept direction for both steering and self-report measurement, testing whether internal-state shifts causally affect the model’s own report of that state. Cross-concept steering uses one concept direction for steering and measures introspection for a different concept, testing whether the modulability of introspective fidelity is concept-specific. This yields a $4\times 4$ matrix of steering-concept $\times$ measured-concept conditions (16 cells $\times$ 5 alphas =80=80 experiments, each over all 40 conversations).
@@ 3.5 Metrics and statistical analysis
Probe-report coupling is assessed with two monotonic association metrics. Spearman $\rho$ measures whether probe scores and self-reports preserve the same ordering across observations, so we refer to it as introspective strength. This metric is positive when self-reports are positively correlated with probe scores (e.g. the model reports higher happiness when the probe measures an internal representation that is closer to happiness) and negative when these variables have an inverse relationship (e.g. the model reports higher happiness when the probe measures an internal representation that is closer to sadness). For this reason, the more positive this metric is in our experiments, the better. If $\rho$ is significantly different from 0, then the association between self-reports and probe scores is not due to chance. Isotonic $R^{2}$ measures how well self-report can be predicted from probe score by an unconstrained monotonic function, so we refer to it as introspective fidelity. In our experiments, higher Isotonic $R^{2}$ can be interpreted as more precise agreement between self-reports and internal state. We report both metrics because a model may preserve rank order on average while the mapping from state to report could be high-variance. Confidence intervals are computed via cluster bootstrap (resampling at the conversation level, B=1, [^0] B=1{,}000, 95% percentile CIs) to respect the non-independence of observations within conversations without relying on parametric normality for rank-based summaries. For pooled conversation-turn and steering observations, significance testing uses linear mixed-effects models (LMM, random intercept by conversation, REML estimation) so that repeated observations from the same conversation do not result in pseudoreplication. Turn-wise and alpha-wise trends are tested with these mixed-effects models when they converge; when the mixed-effects model is singular we fall back to one-sample $t$-tests on per-conversation slopes or per-conversation alpha-correlations, preserving the conversation as the unit of analysis but losing some information. Whenever a figure reports multiple parallel tests from the same dataset and endpoint family, we report raw p-values but apply Benjamini–Hochberg (BH) correction within that family and note whether the reported conditions survive. Comparisons between true probes and random-direction controls use paired cluster-bootstrap tests. Drift across turns is quantified via mixed-effects models (value $\sim$ turn ++ (1||conversation)), reporting the fixed-effect slope as the primary estimate. When only five aggregated alpha means are available, exact-permutation tests over the five alpha levels are used to test monotonic trends. Model-size trends on conversation-level summaries use OLS regression on $\log(\text{model size})$. A significance threshold of $\alpha=0.05$ was used in all analyses.
@@ 4.1 Default self-reports mask internal-state drift
We first asked whether, during ordinary multi-turn conversation, the internal emotive states of the LLM change over time and, if they do, whether the model would report those changes when asked.
@@
We began by asking the model to rate its current state on one of 4 emotive states (wellbeing, interest, focus, and impulsivity) on a 0–9 integer scale and recorded its output, obtained by using greedy decoding over the logit distribution (i.e. we picked the highest probability token). Fig. 2A shows the greedy self-reports across turns for all four concepts. The reports are largely uninformative: the model tends to produce the same number at every turn, especially for the focus and impulsivity concepts (notice that turns 4–10 for focus and 1–8 for impulsivity all have a variance of 0). Interest and wellbeing are more informative, but still collapse to relatively few values. Across concepts, self-reports occupy on average only 1.1–3.9 distinct values out of the 10-point scale (Fig. 2B, more detail on this in Appendix C). Despite this low dynamic range, close inspection reveals that greedy reports are not entirely static over time. Interest drifts upward over the course of a conversation (LMM turn slope =0.14=0.14, $p<10^{-41}$), and smaller positive trends are detectable for wellbeing and focus (slopes =0.029=0.029 and $0.022$, $p<10^{-3}$), though impulsivity does not drift reliably (slope =0.002=0.002, $p=0.20$).
@@
Figure 2: Internal-state drift is tracked by numeric self-reports of the same concept. All panels use 40 ten-turn conversations; shaded bands denote cluster-bootstrap 95% CIs across conversations. Panel A shows greedy integer self-reports across turns, with thin lines for individual conversations and thick lines for per-turn means. Greedy ratings are largely collapsed, with clear positive drift only for interest (mixed-effects turn slope =0.14=0.14, $p=4.98\times 10^{-42}$) and smaller positive trends for wellbeing and focus (slopes =0.029=0.029 and $0.022$, $p=8.38\times 10^{-4}$ and $3.37\times 10^{-6}$); impulsivity does not drift reliably (slope =0.002=0.002, $p=0.20$). Panel B shows the average number of distinct greedy responses used for each concept across the 40 conversations, averaging across turns. Panel C shows probe scores from the trained concept directions across turns. Probe scores drift strongly for interest and focus (slopes =0.005=0.005 and $0.002$, $p=4.12\times 10^{-14}$ and $1.75\times 10^{-10}$), are flat for wellbeing (slope $3.6\times 10^{-4}$, $p=0.64$), and show only a weak effect for impulsivity (LMM slope =0.001=0.001, $p=0.002$). Panel D shows logit-based self-reports, defined as the probability-weighted average over the digit-token logits. Unlike greedy or sampled outputs, the logit-based measure shows robust drift in all four concepts, with positive slopes for wellbeing, interest, and focus and a negative slope for impulsivity (all $p<10^{-6}$). Panel E shows Shannon entropy of the self-report distribution. The logit-based method is most informative in all four concepts. Within each four-concept family in panels A, C, and D, BH correction leaves the significance pattern unchanged.
@@
To determine whether these report-level trends reflect genuine internal-state dynamics, we turned to the trained concept probes (Section 3.3). Fig. 2C shows the probe scores across turns for each of the four concepts. Interest and focus drift reliably towards the positive pole of the concept over the course of conversation (LMM slopes =0.005=0.005 and $0.002$, $p<10^{-9}$, meaning drift towards interest and away from boredom, and towards focused and away from distracted, respectively), while wellbeing remains relatively flat (slope $3.6\times 10^{-4}$, $p=0.64$). Impulsivity is weaker and directionally mixed: its average LMM slope is slightly positive (slope =0.001=0.001, $p=0.002$), but the net first-to-last change is slightly negative, indicating a small non-monotonic trajectory rather than a clean temporal drift. Notably, even though individual conversations (thin lines in Fig. 2C) vary appreciably from one another, they follow a consistent average trend (note in particular the interest and impulsivity panels, where a rapid initial drift is followed by a smaller rebound, and this is coherent across dialogues). This shows that probe directions are capturing a meaningful underlying state change over time, not an artifact of a few outlier conversations.
@@
Given that greedy self-reports are mostly collapsed, we reasoned that it would be desirable to have a more informative metric as a self-reported measure of internal state if one wanted to track these changes. We first attempted to do non-greedy sampling of the logit distribution to get the self-reports (see Appendix C), which helped moderately but still resulted in a relatively collapsed distribution of a few discrete values. Furthermore, this strategy helps only by adding noise to the output, it does not recover information that is inherently lost during the sampling process. Therefore, we next asked whether there is a way to extract richer self-report information from the model.
@@
Token sampling selects output tokens based on the probability distribution produced by the model at each step, but this process discards a lot of information present in that distribution (Kumar et al., 2024; Geng et al., 2023; Yona et al., 2024). In this case, we care about the full distribution over the ten digit tokens (0–9), which can be more informative than the highest probability sample. We therefore computed a logit-based self-report (Eq. 1): instead of taking a single token, we compute the probability-weighted expected value over all ten digit-token logits, yielding a continuous rating that preserves the model’s full distributional signal.
Fig. 2D shows that this measure reveals robust drift in all four concepts: wellbeing, interest, and focus drift positively, while impulsivity drifts negatively (towards planning); (all $p<10^{-6}$; LMM turn slopes). The direction of logit-based drift matches probe drift for interest and focus, the two cases with clean positive probe drift over time, suggesting that logit-based self-reports are capturing information about internal-state changes. For impulsivity, the logit-based self-report drifts negative while the probe shows only a weak non-monotonic temporal pattern. Notably, even when probe drift is not significant (i.e. for the wellbeing probe), this self-reported metric does show significant change over time, which is consistent with the possibility that self-report captures aspects of internal state that are not fully captured by a single linear probe. Fig. 2E quantifies the improvement achieved by using this technique: the Shannon entropy of the self-report distribution is highest for the logit-based method across all four concepts, ranging from 3.1 to 3.7 bits, confirming that it can yield substantially more information than either greedy decoding (0.03 to 1.1 bits) or sampled decoding (0.68 to 2.1 bits). More information about logit-vs-sampled calibration are provided in Appendix C. The logit-based self-report is used as the primary self-report measure in all subsequent analyses.
@@ 4.2 Self-reports correlate with probe-defined internal state
Logit-based self-report capture meaningful variation that tracks internal-state drifts. However, it is unclear so far if the self-report is actually carrying information about the model’s internal state at the level of individual observations (i.e. if the model is more likely to rate its emotive state higher when the matching linear probe yields a higher score). To answer this question, we explored whether the model’s numeric report of a given emotive state covaries monotonically with the corresponding probe-defined internal direction.
@@
Figure 3: Self-reports track the probe-defined internal state from the first turn, and direct self-steering shifts self-reports causally in the predicted direction. Panel A shows probe score versus logit-based self-report, with one point per conversation-turn observation and black isotonic fits. Descriptive associations are positive for all four concepts (pooled $\rho=0.40$$0.76$; isotonic $R^{2}=0.12$$0.54$), and mixed-effects probe slopes are positive in all cases (all $p<10^{-5}$). Panel B shows turn-wise Spearman $\rho$ (introspective strength) for the four concepts; panel C shows turn-wise isotonic $R^{2}$ (introspective fidelity). Shaded bands denote bootstrap 95% CIs. Introspection is already present at turn 1 and remains positive through turn 10, although its trajectory is concept-dependent. Panel D shows mean logit-based self-report versus steering alpha; shaded bands denote cluster-bootstrap 95% CIs. Steering shifts self-report monotonically in the expected direction for all four concepts (mixed-effects alpha slopes 0.0670.0670.400.40, all $p<10^{-12}$). Panel E shows drift magnitude (last minus first turn) versus steering alpha, with cluster-bootstrap 95% CIs. Wellbeing drift becomes slightly more positive with alpha (LMM slope =0.012=0.012, $p=4.30\times 10^{-6}$), impulsivity drift becomes more negative with alpha (LMM slope =0.028=-0.028, $p=8.92\times 10^{-53}$), and interest and focus drift decrease with alpha (fallback per-conversation slope means =0.095=-0.095 and 0.11-0.11; $p=7.45\times 10^{-14}$ and $1.57\times 10^{-14}$, because the corresponding LMMs are singular). Across the four-concept families in panels A, D, and E, all reported effects survive BH correction across concepts.
These correlations could in principle be driven by probe artifacts rather than genuine internal-state information. To control for this, we repeated the analysis replacing each trained probe with an equal-norm random vector. True probes outperform random controls for all four concepts in both metrics ($\Delta\rho=0.23$$0.79$, all $p<0.05$; $\Delta R^{2}=0.09$$0.44$, all $p<0.05$; paired cluster-bootstrap; all four concept-wise comparisons remain significant after BH correction). Examined on their own, the random-direction controls were weak ($\rho=-0.11$ to $0.17$; isotonic $R^{2}=0.03$$0.11$), and none showed a significant monotonic correlation with self-report under the same two-sided cluster-bootstrap test (wellbeing $p=0.328$, interest $p=0.324$, focus $p=0.090$, impulsivity $p=0.795$). This shows that self-reports carry information specifically about the probe-defined concept directions, and don’t show a significant correlation with random directions in latent space.
4.3 Introspection is present from the first turn and evolves over time

### 11 Apr 2026 v2

@@ 3.3 Concept probes
We train probes for four emotive concept pairs: sad vs. happy (wellbeing), bored vs. interested (interest), distracted vs. focused (focus), and impulsive vs. planning (impulsivity). Each probe is a contrastive mean-difference direction (Zou et al., 2023): for a set of between 20 and 24 neutral questions (e.g., “Explain how photosynthesis works”), we generate completions under two opposing system prompts that induce the positive and negative poles of the concept, extract hidden-state activations from all layers, and compute the concept vector at each layer $l$ as:
@@
where $\bar{\mathbf{h}}_{l}^{+}$ and $\bar{\mathbf{h}}_{l}^{-}$ are the mean hidden-state representations (averaged over all assistant tokens) of completions under the positive and negative system prompts, respectively. We select the best layer $l^{*}$ as the one maximizing Cohen’s $d$ on a set of held-out evaluation texts, with the search restricted to the middle 60% of layers (layers 5–23 in the 28-layer 3B model; Appendix B gives the corresponding ranges for the other models), to lower the chance of selecting layers that mainly capture syntactic or grammatical differences between prompt and/or response structure. The probe score window then includes $l^{*}$ and its two nearest neighbours on each side. Full details on system prompts, training questions, evaluation texts, and sign conventions are provided in Appendix B.
@@
Figure 1: Method overview and probe validation in LLaMA-3.2-3B-Instruct. Panel A schematizes the conversational measurement setup: Gemini 2.5 Flash acts as the simulated user, the model under study acts as the assistant, 40 conversations of 10 turns are generated, and at each turn an independent concept-matched 0–9 rating question is appended, yielding one self-report and one previous-turn probe score from the preceding assistant response. Panels B, D, F, and H show layer-wise Cohen’s $d$ sweeps for the sad-vs-happy (wellbeing), bored-vs-interested (interest), distracted-vs-focused (focus), and planning-vs-impulsive (impulsivity) probes; dashed lines mark the selected layers and the gray band marks the searched layer range (layers 5–23). The best layers are 16, 14, 10, and 13, with peak $d$ values 3.34, 1.67, 1.99, and 3.60. For wellbeing and impulsivity, scores were sign-corrected so that larger values align with larger self-reported values in later experiments. Panels C, E, G, and I show the selected-layer score distributions on held-out evaluation texts; boxplots show text-level scores. The two poles separate in all four cases (Welch’s $t$-test: wellbeing, $d=3.34$, $p=7.21\times 10^{-13}$; interest, $d=1.67$, $p=9.45\times 10^{-6}$; focus, $d=1.99$, $p=5.73\times 10^{-7}$; impulsivity, $d=3.60$, $p=3.58\times 10^{-13}$), and all four survive BH correction across concepts.
@@
Figure 1 first summarizes the measurement pipeline and then presents the probe-quality results for all four concepts in LLaMA-3.2-3B-Instruct. Panels B, D, F, and H show the layer-wise Cohen’s $d$ sweep for each probe; the selected best layers are 16, 14, 10, and 13, for wellbeing, interest, focus, and impulsivity respectively. Panels C, E, G, and I show the score distributions on held-out evaluation texts at the selected layer. The two poles separate clearly in all four cases (Welch’s $t$-test: all $p<10^{-5}$; Cohen’s d=1.67d=1.673.603.60), confirming that the trained directions capture meaningful structure. These probes serve as the internal-state readouts for all subsequent experiments.
@@ 3.4 Activation steering
Same-concept steering uses the same concept direction for both steering and self-report measurement, testing whether internal-state shifts causally affect the model’s own report of that state. Cross-concept steering uses one concept direction for steering and measures introspection for a different concept, testing whether the modulability of introspective fidelity is concept-specific. This yields a $4\times 4$ matrix of steering-concept $\times$ measured-concept conditions (16 cells $\times$ 5 alphas =80=80 experiments, each over all 40 conversations).
@@ 3.5 Metrics and statistical analysis
Probe-report coupling is assessed with two monotonic association metrics. Throughout Results, Spearman $\rho$ is reported as introspective strength and isotonic $R^{2}$ as introspective fidelity. We report both because a model may preserve rank order on average while the mapping from state to report could still be high-variance. Confidence intervals are computed via cluster bootstrap (resampling at the conversation level, B=1, [^0] B=1{,}000, 95% percentile CIs) to respect the non-independence of observations within conversations without relying on parametric normality for rank-based summaries.
For pooled conversation-turn observations, significance testing uses linear mixed-effects models (LMMs; random intercept by conversation, REML estimation). The main model equations are: logit_report $\sim$ probe ++ (1||conversation) for pooled introspection, value $\sim$ turn ++ (1||conversation) for turn-wise drift, value $\sim$ alpha ++ (1||conversation) for steering trends, and logit_report $\sim$ probe $\times$ turn ++ (1||conversation) for time-varying coupling. These models keep repeated observations from the same conversation from counting as independent.
Turn-wise and alpha-wise trends are tested with these mixed-effects models when they are estimable; when the LMM is singular we fall back to one-sample $t$-tests on per-conversation slopes or per-conversation alpha-correlations. This fallback preserves the conversation as the unit of analysis (avoids pseudorreplication), but uses less information than the pooled LMM. Whenever a figure reports multiple parallel tests from the same dataset and endpoint family, we report raw p-values but apply Benjamini–Hochberg (BH) correction, a standard false-discovery-rate procedure, within that family and note whether the reported conditions survive. Comparisons between true probes and random-direction controls use paired cluster-bootstrap tests. When only five aggregated alpha means are available, exact-permutation tests over the five alpha levels are used to test monotonic trends. Model-size trends on conversation-level summaries use OLS regression on $\log(\text{model size})$. A significance threshold of $\alpha=0.05$ was used in all analyses.
@@ 4.1 Default self-reports mask internal-state drift
We first asked whether, during multi-turn conversations, the internal emotive states of the LLM change over time and, if they do, whether the model would report those changes when asked.
@@
We began by asking the model to rate its current state on one of 4 emotive states (wellbeing, interest, focus, and impulsivity) on a 0–9 integer scale and recorded its output, obtained by using greedy decoding over the logit distribution (i.e. we picked the highest probability token). Fig. 2A shows the greedy self-reports across turns for all four concepts. The reports are largely uninformative: the model tends to produce the same number at every turn, especially for the focus and impulsivity concepts (notice that turns 4–10 for focus and 1–8 for impulsivity all have a variance of 0). Interest and wellbeing are more informative, but still collapse to relatively few values and often yield nearly flat trajectories. Across concepts, self-reports occupy on average only 1.1–3.9 distinct values out of the 10-point scale (see unique values averaged across turns for each concept in Fig. 2A). Despite this low dynamic range, close inspection reveals that greedy reports are not entirely static over time. Interest drifts upward over the course of a conversation (LMM turn slope =0.14=0.14, $p<10^{-41}$), and smaller positive trends are detectable for wellbeing and focus (slopes =0.029=0.029 and $0.022$, $p<10^{-3}$), though impulsivity does not drift reliably (slope =0.002=0.002, $p=0.20$).
@@
Figure 2: Internal-state drift is tracked by numeric self-reports of the same concept. All panels use 40 ten-turn conversations; shaded bands denote cluster-bootstrap 95% CIs across conversations. Panel A shows greedy integer self-reports across turns, with thin lines for individual conversations and thick lines for per-turn means. Greedy ratings are largely collapsed, using only 1.1–3.9 distinct values on average across conversations (‘Unique’) and entropies of 0.03–1.10 bits (‘Entropy’); only interest shows strong positive drift (mixed-effects turn slope =0.14=0.14, $p=4.98\times 10^{-42}$), while wellbeing and focus show smaller positive trends ($0.029$ and $0.022$; $p=8.38\times 10^{-4}$ and $3.37\times 10^{-6}$) and impulsivity does not drift reliably ($0.002$, $p=0.20$). Panel B shows probe scores from the trained concept directions across turns. Probe scores drift strongly for interest and focus ($0.005$ and $0.002$; $p=4.12\times 10^{-14}$ and $1.75\times 10^{-10}$), are nearly flat for wellbeing ($3.6\times 10^{-4}$, $p=0.64$), and are weakly non-monotonic for impulsivity ($0.001$, $p=0.002$). Panel C shows logit-based self-reports, the probability-weighted average over digit-token logits. This continuous measure has substantially higher entropy (3.1–3.7 bits) and shows robust drift in all four concepts (wellbeing =0.0167=0.0167, interest =0.0988=0.0988, focus =0.0422=0.0422, impulsivity =0.0127=-0.0127; all $p<10^{-6}$). The analogous sampled-decoding analysis is shown in Appendix C. Within each four-concept family in panels A–C, BH correction leaves the significance pattern unchanged.
@@
To determine whether these report-level trends reflect genuine internal-state dynamics, we turned to the trained concept probes (Section 3.3). Fig. 2B shows the probe scores across turns for each of the four concepts. Interest and focus drift reliably towards the positive pole of the concept over the course of conversation (LMM slopes =0.005=0.005 and $0.002$, $p<10^{-9}$, meaning drift towards interest and away from boredom, and towards focused and away from distracted, respectively), while wellbeing remains relatively flat (slope $3.6\times 10^{-4}$, $p=0.64$). Impulsivity is weaker and directionally mixed: its average LMM slope is slightly positive (slope =0.001=0.001, $p=0.002$), but the net first-to-last change is slightly negative, indicating a small non-monotonic trajectory rather than a clean temporal drift. Notably, even though individual conversations (thin lines in Fig. 2B) vary appreciably from one another, they follow a consistent average trend (note in particular the interest and impulsivity panels, where a rapid initial drift is followed by a smaller rebound, and this is coherent across dialogues). This shows that probe directions are capturing a meaningful underlying state change over time, not an artifact of a few outlier conversations.
@@
Given that greedy self-reports are mostly collapsed, we reasoned that it would be desirable to have a more informative metric as a self-reported measure of internal state if one wanted to track these changes. We first attempted non-greedy sampling of the logit distribution (Appendix C, Fig. 6B), which helped moderately but still resulted in a relatively collapsed distribution of a few discrete values. Furthermore, this strategy helps only by adding noise to the output; it does not recover information that is inherently lost during the sampling process. Therefore, we next asked whether there is a way to extract richer self-report information from the model.
@@
Token sampling selects output tokens based on the probability distribution produced by the model at each step, but this process discards a lot of information present in that distribution (Kumar et al., 2024; Geng et al., 2023; Yona et al., 2024). In this case, we care about the full distribution over the ten digit tokens (0–9), which can be more informative than the highest probability sample. Prior work has shown that finer-grained scalar judgments can be extracted from token distributions when LLMs are asked to produce numerical ratings (Zawistowski, 2024; Wang et al., 2025b). We therefore computed a logit-based self-report (Eq. 1): instead of taking a single token, we compute the probability-weighted expected value over all ten digit-token logits, yielding a continuous rating that preserves the model’s full distributional signal.
Fig. 2C shows that this measure reveals robust drift in all four concepts: wellbeing, interest, and focus drift positively, while impulsivity drifts negatively (towards planning); the corresponding LMM turn slopes are $0.0167$, $0.0988$, $0.0422$, and 0.0127-0.0127, respectively (Table 1). The direction of logit-based drift matches probe drift for interest and focus, the two cases with clean positive probe drift over time, suggesting that logit-based self-reports are capturing information about internal-state changes. For impulsivity, the logit-based self-report drifts negative while the probe shows only a weak non-monotonic temporal pattern. Notably, even when probe drift is not significant (i.e. for the wellbeing probe), this self-reported metric does show significant change over time, which is consistent with the possibility that self-report captures aspects of internal state that are not fully captured by a single linear probe. To quantify the improvement achieved by using this technique, we measured the Shannon entropy of the self-report distribution (’Entropy’ values for Fig. 2A and Fig. 2C): logit-based self-report has the highest entropy in all four concepts, ranging from 3.1 to 3.7 bits, versus 0.03 to 1.10 bits for greedy decoding and 0.68 to 2.05 bits for sampled decoding (see Appendix C for the full sampled and calibration analyses). The logit-based self-report is used as the primary self-report measure in all subsequent analyses.
@@ 4.2 Self-reports correlate with probe-defined internal state
Logit-based self-report captures meaningful variation that tracks internal-state drifts. However, it is unclear so far if the self-report is actually carrying information about the model’s internal state at the level of individual observations (i.e. if the model is more likely to rate its emotive state higher when the matching linear probe yields a higher score). To answer this question, we explored whether the model’s numeric report of a given emotive state covaries monotonically with the corresponding probe-defined internal direction. Following the definitions in Section 3.2, we summarize this coupling with two metrics: introspective strength (Spearman $\rho$) and introspective fidelity (isotonic $R^{2}$). If $\rho$ is significantly different from 0, then the association between self-reports and probe scores is not due to chance. Furthermore, higher introspective fidelity can be interpreted as more precise agreement between self-reports and internal state.
@@
Figure 3: Self-reports track the probe-defined internal state from the first turn, and introspective coupling evolves through conversation. Panel A shows probe score versus logit-based self-report, with one point per conversation-turn observation and black isotonic fits. Descriptive associations are positive for all four concepts (pooled Spearman $\rho=0.40$$0.76$; isotonic $R^{2}=0.12$$0.54$), and mixed-effects probe slopes are positive in all cases (all $p<10^{-5}$). Panel B combines turn-wise introspective fidelity (solid lines; isotonic $R^{2}$) and introspective strength (dashed lines; Spearman $\rho$) for the four concepts, with bootstrap 95% CIs. Introspection is already present at turn 1 and remains positive through turn 10, although its trajectory is concept-dependent.
These correlations could in principle be driven by probe artifacts rather than genuine internal-state information. To control for this, we repeated the analysis replacing each trained probe with an equal-norm random vector: one isotropic Gaussian direction per layer, normalized and then rescaled to match the trained vector norm at that layer. True probes outperform random controls for all four concepts in both metrics ($\Delta\rho=0.23$$0.79$, all $p<0.05$; $\Delta R^{2}=0.09$$0.44$, all $p<0.05$; paired cluster-bootstrap; all four concept-wise comparisons remain significant after BH correction). Examined on their own, the random-direction controls were weak ($\rho=-0.11$ to $0.17$; isotonic $R^{2}=0.03$$0.11$), and none showed a significant monotonic correlation with self-report under the same two-sided cluster-bootstrap test (wellbeing $p=0.328$, interest $p=0.324$, focus $p=0.090$, impulsivity $p=0.795$). This shows that self-reports carry information specifically about the probe-defined concept directions, and don’t show a significant correlation with random directions in latent space.
4.2.1 Introspection is present from the first turn and evolves over time